Solve the equations by substitution.
8x = 2y + 5
3x = y + 7
x = ?
y = ?

Found 2 solutions by checkley77, palanisamy:
Answer by checkley77(12844)

8x = 2y + 5
3x = y + 7 OR Y=3X-7
8X=2(3X-7)+5
8X=6X-14+5
8X-6X=-9
2X=-9
X=-9/2
X=-4.5 ANS.
8*-4.5=2Y+5
-36=2Y+5
2Y=-36-5
2Y=-41
Y=-41/2
Y=-20.5 ANS.
PROOF:
3*-4.5=-20.5+7
-13.5=-13.5


Answer by palanisamy(496)

The given equations are,
8x = 2y + 5 ...(1)
3x = y + 7
3x-7 = y
y = 3x-7
Substituting in (1), we get
3x = 2(3x-7)+5
3x = 6x-14+5
3x-6x = -9
-3x = -9
x = (-9)/(-3)
x = 3
Substituting in (1), we get
3*(3) = 2y+5
9-5 = 2y
2y = 4
y = 4/2
y = 2
So the solution is x = 3, y = 2